# Bridge multisections of knotted surfaces in $S^{4}$

### v1

@@ 6. Band surgeries are unknotting moves
Let $N(\Sigma)$ be a small tubular neighborhood of $\Sigma$ in $S^{4}$. We choose $N(\Sigma)$ so that $N(\Sigma)\cap F$ is the disjoint union of $2b$ disks around the bridge points. For each $i=1,\dots,n$, denote by $N_{i}(\Sigma)=N(\Sigma)\cap B_{i}$ the a small neighborhood of $\Sigma$ in $B_{i}$ with boundary $\Sigma\cup\Sigma_{i}$. Note that $\Sigma_{i}$ intersects $T_{i}$ in $2b$ points and the subset of $T_{i}$ inside $B_{i}\setminus int(N(\Sigma))$ is still a $b$–bridge trivial tangle. By construction, the spheres $\Sigma_{1},\dots,\Sigma_{n}$ are a subset of $\partial N(\Sigma)\approx\Sigma\times S^{1}$.
@@
To end, suppose that $j=2$ (the same argument works for $j=n$). The tuple of arcs $\rho$ defined in the previous paragraph may not be admissible. To ensure the admissibility condition, we perform extra multiple-sector perturbations described in Figure 22. First, we use the increasing arcs $t\cap B_{1}$ and $t\cap B_{2}$ to guide two consecutive 0–sector perturbations. We follow this with consecutive 1–sector perturbations as in Figure 22. The link $T_{1}\cup\overline{T}_{2}$ for the resulting multiplane diagram has two new 1–bridge unknots (obtained from the 1–sector perturbations). Moreover, the guiding arc $t$ can be slid to connect those two new unknots as in Figure 22. In particular, we obtain an admissible tuple of $\rho$ with $\rho_{1}=t\cap B_{1}$, $\rho_{2}=t\cap B_{2}$, and $\rho_{i}$ type 0 for $i\neq 1,2$. By Lemma 4.3, $(\rho,f)$ is an admissible tuple for any framing $f$. Furthermore, by Example 4.6, banded surgery along $(\rho,f)$ represents a 1–handle addition to $F$ with guiding arcs $\rho_{1}\cup\rho_{2}=t$. This finishes the proof of the proposition. ∎
@@
We are ready to prove the main results of this section. This relies heavily on [HK79, Thm 2.3] which states that any surface in 4-space can be unknotted by adding enough 1–handles. In particular, any pair of surfaces can be related by a sequence of 1–handle additions, connected sums with unknotted projective spaces, and their inverses.
@@
Let $\mathcal{T}$ be an arbitrary multiplane diagram. After a finite number of band surgeries, the resulting multiplane diagram $\mathcal{T}^{\prime}$ satisfies that $F_{\mathcal{T}^{\prime}}$ is an unknotted surface.
@@
By [HK79, Thm 2.3], we know that any surface in 4-space can be unknotted by adding enough 1–handles. Proposition 6.1 ensures that each 1–handle addition can be performed via finitely many band surgeries to $\mathcal{T}$. ∎
@@
Let $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ be two multiplane diagrams representing the surfaces $F_{1}$ and $F_{2}$, respectively. By [HK79, Thm 2.3], there is a sequence of 1–handle additions, crosscap sums, and their inverses, taking $F_{1}$ into $F_{2}$. By Proposition 6.1 and Example 4.8, arbitrary 1–handle additions and crosscap sums can be performed using band surgeries on multiplane diagrams. Thus, we can find a finite sequence of band surgeries and their inverses, that turn $\mathcal{T}_{1}$ into a multiplane diagram $\mathcal{T}^{\prime}_{1}$ representing the surface $F_{2}$. Hence, by Theorem 5.8 we contain a finite sequence of multiplane moves turning $\mathcal{T}^{\prime}_{1}$ into $\mathcal{T}_{2}$. ∎
@@ 7. Graphs induced by bridge multisections
An edge coloring of a graph is a proper coloring of the edges, meaning an assignment of colors to edges so that no vertex is incident to two edges of the same color. An edge coloring with $n$ colors is called an $n$–edge coloring. The spine of a bridge multisection is an $n$–valent graph with an edge coloring where the edges of color $i$ are the arcs of the $i$-th tangle $T_{i}$. This section aims to show that all $n$–edge colored $n$–valent graphs appear this way. We do this by studying how the graph type of the multisection spine changes under band surgery along an admissible tuple $(\rho,f)$.
Let $\Gamma$ be an $n$–valent graph with an $n$–edge coloring. Given an edge $e$, we can modify $\Gamma$ by an operation similar to an edge-contraction that preserves the given $n$–edge coloring. We define the edge-compression of $\Gamma$ along $e$, denoted by $\Gamma|e$, as follows: The vertex set of $\Gamma|e$ equals $V(\Gamma)\setminus\partial e$. Edges parallel to $e$ in $G$ get deleted in $G|e$. There are two types of edges for $\Gamma|e$. If an edge $f$ in $\Gamma$ is not adjacent nor parallel to $e$, then $f$ is also an edge of $\Gamma|e$ with the same color. If two distinct edges $f$ and $g$ of $\Gamma$ are adjacent to $e$ and have the same color, then we add a new edge to $\Gamma|e$ that connects the endpoints of $f$ and $g$ not in $e$ with the same color as $f$ and $g$. Figure 23 has examples of edge-compressions. By construction, $\Gamma|e$ is also an $n$–valent graph with an $n$–edge coloring that agrees in the common edges of $\Gamma$ and $\Gamma|e$.
@@
As seen in [MTZ23, Sec 3.5], at the level of the 1-skeleta, edge compressions are inverses of band surgeries along admissible tuples of arcs. More precisely, consider two multiplane diagrams $\mathcal{T}_{0}$ and $\mathcal{T}_{1}$ related by band surgery along some admissible tuple of arcs $(\rho,f)$; i.e. $\mathcal{T}_{1}=\mathcal{T}_{0}[\rho,f]$. If $\rho$ has at least one type 0 arc, then the skeleton of $\mathcal{T}_{0}$ is isomorphic to an edge-compression on the skeleton of $\mathcal{T}_{1}$. The following lemma shows that this relation can be reversed.
@@
Let $\Gamma_{1}$ be an $n$–valent graph with an $n$–edge coloring. Assume that there is an edge $e$ of $\Gamma_{1}$ and a multiplane diagram $\mathcal{T}_{0}$ satisfying that the spine of $\mathcal{T}_{0}$ with its induced $n$–edge color is isomorphic to $\Gamma_{1}|e$. Then there exists an admissible tuple $(\rho,f)$ for $\mathcal{T}_{0}$ satisfying that the spine of $\mathcal{T}_{0}[\rho,f]$ with its induced coloring is isomorphic to $\Gamma_{1}$.
@@
In what follows, we will recursively build $\{(\rho_{i+1},f_{i+1})\}_{i=1}^{n}$. Without loss of generality, we assume that the color of $e$ is $1$. This way, $(\rho_{1},f_{1})$ can be chosen to be any type 0 framed arc in $T_{1}$. Note that, regardless of our final choice for $(\rho_{n},f_{n})$, the pair $(\rho_{n},\rho_{1})$ will satisfy the admissibility conditions of Definition 4.1(3-4) as $\rho_{1}$ is already type 0. Assume that the arc $(\rho_{i},f_{i})$ has been defined for some $i=1,\dots,n-1$. In order to choose $(\rho_{i+1},f_{i+1})$, we look at the $(i+1)$–colored edges in $\Gamma$ that are adjacent to $e$. We have two cases, depending on how many such edges are. If there is only one such edge, this edge is parallel to $e$, and we can then choose $\rho_{i+1}$ to be a type 0 arc in $T_{i+1}$ with any framing $f_{i+1}$.
Suppose there exist two $(i+1)$–colored edges in $\Gamma$ adjacent to $e$ denoted by $g$ and $h$. By definition of $\Gamma|e$, the edges $g$ and $h$ get traded with a new $(i+1)$–colored edge connecting the vertices in $(\partial g\cup\partial h)\setminus\partial e$. As $\Gamma|e$ is the spine of $\mathcal{T}_{0}$, this new edge corresponds to a strand $t_{i+1}$ of $T_{i+1}$. If $\rho_{i}$ is type 0, we can choose $\rho_{i+1}$ to be any type 1 arc with one endpoint in $t_{i+1}$ and the admissibility conditions of Definition 4.1(3-4) will be satisfied for $\rho_{i}\cup\rho_{i+1}$. The framing of $\rho_{i+1}$ will be chosen so that $T_{i+1}[\rho_{i+1},f_{i+1}]$ connects the correct endpoints of $g$ and $h$.
We are left to discuss the case when $\rho_{i}$ is a type 1 arc. The following argument takes place in a given diagram for the bridge splitting $(S^{3},L_{i})=(B_{i},T_{i})\cup_{\Sigma}(B_{i+1},T_{i+1})$. First, slide the endpoints of $\rho_{i}$ so that this arc is inside a small neighborhood of the bridge sphere, $N(\Sigma)$. By a result of Solanki [Sol23, Cor 2], there is a sequence of braid isotopies, pocket moves, flip moves, and bridge destabilizations taking $L_{i}=T_{i}\cup T_{i+1}$ to a crossingless diagram of $L_{i}$ in with $|L_{i}|$ bridges. See [Sol23, Sec 2.2] for detailed definitions of these moves. We keep track of $\rho_{i}$ through this sequence and make sure that $\rho_{i}$ is still a decreasing arc lying in $N(\Sigma)$. Note that braid isotopies and flip moves preserve the desired properties of $\rho_{i}$ as they preserve the regular levels of the Morse function associated with the genus zero Heegaard splitting of $S^{3}$. Pocket moves can be pre- and post-composed with braid isotopies so that they are supported away from $N(\Sigma)$, thus preserving the decreasing property of $\rho_{i}$. Every time we find a destabilization move in our sequence, instead of canceling the pair of critical points, we can push the $(\min,\max)$ pair to be destabilized to lie inside $N(\Sigma)$ as in Figure 24. Then, future moves in our sequence can be replicated for this $b$–bridge presentation.
@@
At the end of the above process, we are left with a $b$–bridge crossingless diagram for $L_{i}$ and a type 1 arc $\rho_{i}$ that braids with some strands of $L_{i}$. We can then choose $\rho_{i+1}$ so that it connects {}$\{*\}=\rho_{i}\cap int(\Sigma)$ with our strand $t_{i+1}$ while undoing all the braiding caused by $\rho_{i}$. This way, $\rho_{i}\cup\rho_{i+1}$ can be isotoped (via Reidemeister II moves) to have no crossings with $L_{i}$. Thus, $\rho_{i}\cup\rho_{i+1}$ satisfies the admissibility conditions in Definition 4.1(3-4). To end, we pick the framing $f_{i+1}$ so that $f_{i}+f_{i+1}$ is zero or ±$\pm 1$, depending on which endpoint of $e$ (puncture of $\Sigma$) is connected by the arcs of $t_{i+1}[\rho_{i+1},f_{i+1}]$. This concludes the construction of $(\rho,f)$. ∎
@@
Figure 25. Guided by the compressions in Figure 23, we found admissible tuples of arcs turning the 1–bridge 5–section of an unknotted 2–sphere into a 3–bridge 5–section with spine isomorphic to $K_{6}$. One can check how the band surgeries change the multisected surface using the reductions on Remark 4.10.
@@
Let $\Gamma$ be an $n$–valent graph with an $n$–edge coloring. There is a bridge multisection $\mathcal{T}$ for an unknotted surface in $S^{4}$ with 1-skeleton isomorphic to $\Gamma$ and inducing the given edge coloring of $\Gamma$.
@@
We only discuss the case when $\Gamma$ is connected. The disconnected case will follow since the spine of the distant sum of bridge multisections is the disjoint union of their respective spines (see Section 3.3). Let $\{e_{1},\dots,e_{b}\}$ be the edges of $\Gamma$ with color $1$. Recursively, define $\Gamma_{1}=\Gamma$ and $\Gamma_{k+1}=\Gamma_{k}|e_{k}$ for $k=1,\dots,b-1$. By construction, $\Gamma_{b}$ has two vertices and $n$ parallel edges of distinct colors. Thus $\Gamma_{b}$ is the spine of a 1–bridge $n$–section of an unknotted 2–sphere. Denote this multiplane diagram by $\mathcal{T}_{b}$. We now apply Lemma 7.1 multiple times to get a sequence of admissible tuples $(\rho^{1},f^{1}),\dots,(\rho^{b-1},f^{b-1})$ undoing the edge-compressions $\Gamma_{1}\rightarrow\cdots\rightarrow\Gamma_{b}$. More precisely, they satisfy that
$\mathcal{T}_{b-j}=\mathcal{T}_{b-j+1}[\rho^{j},f^{j}]\quad\text{and}\quad\Gamma_{b-j}\text{ is the spine of }\mathcal{T}_{b-j},$

### v2

@@ 6. Band surgeries are unknotting moves
Let $N(\Sigma)$ be a small tubular neighborhood of $\Sigma$ in $S^{4}$. We choose $N(\Sigma)$ so that $N(\Sigma)\cap F$ is the disjoint union of $2b$ disks around the bridge points. For each $i=1,\dots,n$, denote by $N_{i}(\Sigma)=N(\Sigma)\cap B_{i}$ the a small neighborhood of $\Sigma$ in $B_{i}$ with boundary $\Sigma\cup\Sigma_{i}$. Note that $\Sigma_{i}$ intersects $T_{i}$ in $2b$ points and the subset of $T_{i}$ inside $B_{i}\setminus int(N(\Sigma))$ is still a bb–bridge trivial tangle. By construction, the spheres $\Sigma_{1},\dots,\Sigma_{n}$ are a subset of $\partial N(\Sigma)\approx\Sigma\times S^{1}$.
@@
To end, suppose that $j=2$ (the same argument works for $j=n$). The tuple of arcs $\rho$ defined in the previous paragraph may not be admissible, or $\rho_{1}\cup\rho_{2}$ may connect the same component of $T_{1}\cup\overline{T}_{2}$. To ensure the admissibility condition, we perform additional multiple-sector perturbations described in Figure 22. First, we use the increasing arcs $t\cap B_{1}$ and $t\cap B_{2}$ to guide two consecutive 0–sector perturbations. We follow this with consecutive 1–sector perturbations as in Figure 22. The link $T_{1}\cup\overline{T}_{2}$ for the resulting multiplane diagram has two new 1–bridge unknots (obtained from the 1–sector perturbations). Moreover, the guiding arc $t$ can be slid to connect those two new unknots as in Figure 22. In particular, we obtain an admissible tuple of $\rho$ with $\rho_{1}=t\cap B_{1}$, $\rho_{2}=t\cap B_{2}$, and $\rho_{i}$ type 0 for $i\neq 1,2$. Since $\rho_{1}$ and $\rho_{2}$ now connect different components of the link $T_{1}\cup\overline{T}_{2}$, any choice of framing for $\rho_{1}$ and $\rho_{2}$ will make the tuple $(\rho,f)$ admissible. By Example 4.7, banded surgery along $(\rho,f)$ represents a 1–handle addition to $F$ with guiding arcs $\rho_{1}\cup\rho_{2}=t$. This finishes the proof of the proposition. ∎
@@
We are ready to prove the main results of this section. This relies heavily on [HK79, Thm 2.3] which states that any surface in $S^{4}$ can be unknotted by adding enough 1–handles. In particular, any pair of surfaces can be related by a sequence of 1–handle additions, connected sums with unknotted projective spaces, and their inverses.
@@
By [HK79, Thm 2.3], we know that any surface in $S^{4}$ can be unknotted by adding enough 1–handles. Proposition 6.1 ensures that each 1–handle addition can be performed via finitely many band surgeries to $\mathcal{T}$. ∎
@@
Let $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ be two multiplane diagrams representing the surfaces $F_{1}$ and $F_{2}$, respectively. By [HK79, Thm 2.3], there is a sequence of 1–handle additions, crosscap sums, and their inverses, taking $F_{1}$ into $F_{2}$. By Proposition 6.1 and Example 4.9, arbitrary 1–handle additions and crosscap sums can be performed using band surgeries on multiplane diagrams. Thus, we can find a finite sequence of band surgeries and their inverses, that turn $\mathcal{T}_{1}$ into a multiplane diagram $\mathcal{T}^{\prime}_{1}$ representing the surface $F_{2}$. Hence, by Theorem 5.10 there exists a finite sequence of multiplane moves turning $\mathcal{T}^{\prime}_{1}$ into $\mathcal{T}_{2}$. ∎
@@ 7. Graphs induced by bridge multisections
An edge coloring of a graph is a proper coloring of the edges, meaning an assignment of colors to edges so that no vertex is incident to two edges of the same color. An edge coloring with $n$ colors is called an nn–edge coloring. The spine of a bridge multisection is an nn–valent graph with an edge coloring where the edges of color $i$ are the arcs of the $i$-th tangle $T_{i}$. This section aims to show that all nn–edge colored nn–valent graphs appear this way. We do this by studying how the graph type of the multisection spine changes under band surgery along an admissible tuple $(\rho,f)$.
Let $\Gamma$ be an nn–valent graph with an nn–edge coloring. Given an edge $e$, we can modify $\Gamma$ by an operation similar to an edge-contraction that preserves the given nn–edge coloring. We define the edge-compression of $\Gamma$ along $e$, denoted by $\Gamma|e$, as follows: The vertex set of $\Gamma|e$ equals $V(\Gamma)\setminus\partial e$. Edges parallel to $e$ in $G$ get deleted in $G|e$. There are two types of edges for $\Gamma|e$. If an edge $f$ in $\Gamma$ is not adjacent nor parallel to $e$, then $f$ is also an edge of $\Gamma|e$ with the same color. If two distinct edges $f$ and $g$ of $\Gamma$ are adjacent to $e$ and have the same color, then we add a new edge to $\Gamma|e$ that connects the endpoints of $f$ and $g$ not in $e$ with the same color as $f$ and $g$. Figure 23 has examples of edge-compressions. By construction, $\Gamma|e$ is also an nn–valent graph with an nn–edge coloring that agrees in the common edges of $\Gamma$ and $\Gamma|e$.
@@
As seen in [MTZ23, Sec 3.5], at the level of the 1–skeleta, edge compressions are inverses of band surgeries along admissible tuples of arcs. More precisely, consider two multiplane diagrams $\mathcal{T}_{0}$ and $\mathcal{T}_{1}$ related by band surgery along some admissible tuple of arcs $(\rho,f)$; i.e. $\mathcal{T}_{1}=\mathcal{T}_{0}[\rho,f]$. If $\rho$ has at least one type 0 arc, then the skeleton of $\mathcal{T}_{0}$ is isomorphic to an edge-compression on the skeleton of $\mathcal{T}_{1}$. The following lemma shows that this relation can be reversed.
@@
Let $\Gamma_{1}$ be an nn–valent graph with an nn–edge coloring. Assume that there is an edge $e$ of $\Gamma_{1}$ and a multiplane diagram $\mathcal{T}_{0}$ satisfying that the spine of $\mathcal{T}_{0}$ with its induced nn–edge color is isomorphic to $\Gamma_{1}|e$. Then there exists an admissible tuple $(\rho,f)$ for $\mathcal{T}_{0}$ satisfying that the spine of $\mathcal{T}_{0}[\rho,f]$ with its induced coloring is isomorphic to $\Gamma_{1}$.
@@
In what follows, we will recursively build $\{(\rho_{i+1},f_{i+1})\}_{i=1}^{n}$. Without loss of generality, we assume that the color of $e$ is $1$. This way, $(\rho_{1},f_{1})$ can be chosen to be any type 0 framed arc in $T_{1}$. Note that, regardless of our final choice for $(\rho_{n},f_{n})$, the pair $(\rho_{n},\rho_{1})$ will satisfy the admissibility conditions of Definition 4.1(3-4) as $\rho_{1}$ is already type 0. Assume that the arc $(\rho_{i},f_{i})$ has been defined for some $i=1,\dots,n-1$. In order to choose $(\rho_{i+1},f_{i+1})$, we look at the (i+1)(i+1)–colored edges in $\Gamma$ that are adjacent to $e$. We have two cases, depending on how many such edges are. If there is only one such edge, this edge is parallel to $e$, and we can then choose $\rho_{i+1}$ to be a type 0 arc in $T_{i+1}$ with any framing $f_{i+1}$.
Suppose there exist two (i+1)(i+1)–colored edges in $\Gamma$ adjacent to $e$ denoted by $g$ and $h$. By definition of $\Gamma|e$, the edges $g$ and $h$ get traded with a new (i+1)(i+1)–colored edge connecting the vertices in $(\partial g\cup\partial h)\setminus\partial e$. As $\Gamma|e$ is the spine of $\mathcal{T}_{0}$, this new edge corresponds to a strand $t_{i+1}$ of $T_{i+1}$. If $\rho_{i}$ is type 0, we can choose $\rho_{i+1}$ to be any type 1 arc with one endpoint in $t_{i+1}$ and the admissibility conditions of Definition 4.1(3-4) will be satisfied for $\rho_{i}\cup\rho_{i+1}$. The framing of $\rho_{i+1}$ will be chosen so that $T_{i+1}[\rho_{i+1},f_{i+1}]$ connects the correct endpoints of $g$ and $h$.
We are left to discuss the case when $\rho_{i}$ is a type 1 arc. The following argument takes place in a given diagram for the bridge splitting $(S^{3},L_{i})=(B_{i},T_{i})\cup_{\Sigma}(B_{i+1},T_{i+1})$. First, slide the endpoints of $\rho_{i}$ so that this arc is inside a small neighborhood of the bridge sphere, $N(\Sigma)$. By a result of Solanki [SOL23, Cor 2], there is a sequence of braid isotopies, pocket moves, flip moves, and bridge destabilizations taking $L_{i}=T_{i}\cup T_{i+1}$ to a crossingless diagram of $L_{i}$ in with $|L_{i}|$ bridges. See [SOL23, Sec 2.2] for detailed definitions of these moves. We keep track of $\rho_{i}$ through this sequence and make sure that $\rho_{i}$ is still a decreasing arc lying in $N(\Sigma)$. Note that braid isotopies and flip moves preserve the desired properties of $\rho_{i}$ as they preserve the regular levels of the Morse function associated with the genus zero Heegaard splitting of $S^{3}$. Pocket moves can be pre- and post-composed with braid isotopies so that they are supported away from $N(\Sigma)$, thus preserving the decreasing property of $\rho_{i}$. Every time we find a destabilization move in our sequence, instead of canceling the pair of critical points, we can push the $(\min,\max)$ pair to be destabilized to lie inside $N(\Sigma)$ as in Figure 24. Then, future moves in our sequence can be replicated for this bb–bridge presentation.
@@
At the end of the above process, we are left with a bb–bridge crossingless diagram for $L_{i}$ and a type 1 arc $\rho_{i}$ that braids with some strands of $L_{i}$. We can then choose $\rho_{i+1}$ so that it connects $\{*\}=\rho_{i}\cap int(\Sigma)$ with our strand $t_{i+1}$ while undoing all the braiding caused by $\rho_{i}$. This way, $\rho_{i}\cup\rho_{i+1}$ can be isotoped (via Reidemeister II moves) to have no crossings with $L_{i}$. Thus, $\rho_{i}\cup\rho_{i+1}$ satisfies the admissibility conditions in Definition 4.1(3-4). To end, we pick the framing $f_{i+1}$ so that $f_{i}+f_{i+1}$ is zero or $\pm 1$, depending on which endpoint of $e$ (puncture of $\Sigma$) is connected by the arcs of $t_{i+1}[\rho_{i+1},f_{i+1}]$. This concludes the construction of $(\rho,f)$. ∎
@@
Figure 25. Guided by the compressions in Figure 23, we found admissible tuples of arcs turning the 1–bridge 5–section of an unknotted 2–sphere into a 3–bridge 5–section with spine isomorphic to $K_{6}$. One can check how the band surgeries change the multisected surface using the reduction rules on Remark 4.11; see Example 4.15.
@@
Let $\Gamma$ be an nn–valent graph with an nn–edge coloring. There is a bridge multisection $\mathcal{T}$ for an unknotted surface in $S^{4}$ with 1–skeleton isomorphic to $\Gamma$ and inducing the given edge coloring of $\Gamma$.
@@
We only discuss the case when $\Gamma$ is connected. The disconnected case will follow since the spine of the distant sum of bridge multisections is the disjoint union of their respective spines (see Section 3.3). Let $\{e_{1},\dots,e_{b}\}$ be the edges of $\Gamma$ with color $1$. Recursively, define $\Gamma_{1}=\Gamma$ and $\Gamma_{k+1}=\Gamma_{k}|e_{k}$ for $k=1,\dots,b-1$. By construction, $\Gamma_{b}$ has two vertices and $n$ parallel edges of distinct colors. Thus $\Gamma_{b}$ is the spine of a 1–bridge nn–section of an unknotted 2–sphere. Denote this multiplane diagram by $\mathcal{T}_{b}$. We now apply Lemma 7.1 multiple times to get a sequence of admissible tuples $(\rho^{1},f^{1}),\dots,(\rho^{b-1},f^{b-1})$ undoing the edge-compressions $\Gamma_{1}\rightarrow\cdots\rightarrow\Gamma_{b}$. More precisely, they satisfy that
[…]
@@
Remark 7.4.
At this point, the invested reader may wonder if Theorem 7.2 follows from its $n=3$ version proven in [MTZ23, Thm 1.1]; that every 3–valent graph with a 3–edge coloring is the spine of an unlink of unknotted surfaces. Below, we describe an inductive approach one may try that fails. Thus, this motivates the use of our technology in Lemma 7.1. That said, there may still be an alternative proof of Theorem 7.2 unknown to the authors of this work.
First, observe that removing all edges with a fixed color from a $4$-colored graph yields a $3$-valent graph with a $3$-edge coloring. So, by [MTZ23], every 4-colored graph is the spine of a bridge 4-section $\mathcal{T}=(T_{1},T_{2},T_{3},T_{4})$ for $F$ where $\mathcal{T}^{\prime}=(T_{1},T_{2},T_{3})$ and $\mathcal{T}^{\prime\prime}=(T_{3},T_{4},T_{1})$ are spines for unlinks for unknotted surfaces $F^{\prime}$ and $F^{\prime\prime}$, respectively. It is not hard to see that $F$ is built from $F^{\prime}$ and $F^{\prime\prime}$ by removing disks bounded by $T_{3}\cup\overline{T}_{1}$ from both $F^{\prime}$ and $F^{\prime\prime}$ and gluing the resulting surfaces. In other words, $F$ is obtained from $F\sqcup F^{\prime\prime}$ by adding $|T_{1}\cup\overline{T}_{3}|$ 1–handles between the two surface-links. As both $F^{\prime}$ and $F^{\prime\prime}$ are unknotted, one may want to show that $F$ is also unknotted. This is not always the case; e.g., the spun trefoil can be described by one 1–handle addition to the distant sum of two unknotted spheres as in Figure 2. Another comment is that in the same figure, the tuples of tangles (blue, green, purple) and (purple, red, blue) are triplane diagrams for unlinks of 2–spheres while the 4–sected surface is knotted.